# Deep Submillimeter and Radio Observations in the SSA22 Field. III. Multiwavelength Identifications and Properties of the 850 -selected Submillimeter Galaxies

### v1

@@ V.2 SFRs and Stellar Mass
The stellar mass distribution of SMGs in the SSA22 field is consistent with previous studies, where median masses typically range from – M⊙ (cunha2015; Michałowski et al., 2017; Zavala et al., 2018; Dudzevičiūtė et al., 2020b; Shim et al., 2022; Hyun et al., 2023). Although the values are similar in magnitude, the median SFR in our SSA22 sample is 0.1–0.2 dex lower than in the literature. We attribute this to a low-redshift bias in source identification, as high-redshift SMGs—often characterized by higher SFRs and infrared luminosities—may be missed. When we restrict the S2TDF and AS2UDS samples from the literature to , their median SFRs are 220 M, which is comparable to our results.
@@
The median deblended 850 fluxes for counterparts of newly detected sources ( 2.8-3.5 mJy) and S2CLS-detected sources ( 3.5–4.2 mJy) are 3.2 mJy and 4.9 mJy, respectively. These two SMG groups show median SFR and Mstar values of 117 M⊙ yr-1, 1.59 1011 M⊙ and 186 M⊙ yr-1, 1.55 1011 M⊙, correspondingly. These results reveal systematic differences in SFR distributions across flux bins, the relation was derived by least-squares fitting: log10(SFR) = (0.49 0.21) log10(S) + (1.78 0.13). When restricting the sample to sources brighter than 4 (or 5) mJy, the median SFR increases to 192 (296) M⊙ yr-1. However, lower SNR sources do not significantly impact the estimated physical properties. For instance, sources with SNR exhibit median SFR and of 175 M⊙ yr-1 and M⊙, while lower-SNR sources (3.5-4) show values of 130 M⊙ yr-1, 1.39 1011 M⊙.
We adopt the main sequence relation from Speagle et al. (2014). We classify the SMGs exhibiting star formation rates three times above this relation as starburst systems, while those with values three times below are categorized as quenched systems (see Paper IV). SMGs located within the main sequence have median SFR and values of 176 M⊙ yr-1 and M⊙, respectively. The ‘starburst system’ exhbit higher SFR but lower galaxy masses, with median values of SFR 510 M⊙ yr-1 and M⊙. In contrast, the ‘quench system’ are on average more massive, with a median of M⊙ — approximately twice the average of full sample — but a median SFR of only 28 M⊙ yr-1.
@@ V.2.1 Mass-to-light Ratio
At typical redshifts for SMGs, the IRAC 5.8 band corresponds approximately to the rest-frame H-band, which serves as a reliable tracer of the dominant stellar population—low-mass stars (Dudzevičiūtė et al., 2020b). As shown in Figure 18, the 151 SMGs with 5.8 detections follow a distribution and trend consistent with the AS2UDS sample: brighter 5.8 emission corresponds to higher stellar mass. A linear fit yields: [M.
@@ V.3 Infrared Luminosity
The median infrared luminosity (, integrated over 8–1000 in the rest frame) of the sample is L⊙, with a 1 range of – L⊙, and the most luminous source reaches L⊙ (lower-left panel of Figure 17). Similar to the SFR results discussed earlier, while slightly lower than literature values, the infrared luminosities of SSA22 SMGs remain comparable in order of magnitude to previous studies.
@@
Due to the tight correlation between infrared luminosity and star formation activity, exhibits a significant evolvement with the cosmic star formation history in Figure 17. Dudzevičiūtė et al. (2020b) found that over the redshift range –4, scales approximately as , but they argue this trend may be biased by the dependence of SED fitting on FIR data. Nevertheless, it is evident that the most luminous SMGs show a clear decline in brightness: HLIRGs are predominantly found at , and at , SMG infrared luminosities typically fall below L⊙.
@@
As reported by cunha2015; Dudzevičiūtė et al. (2020b), MAGPHYS allow older stellar populations heat dust, contributing excess luminosity—particularly prominent in galaxies with higher mass-weighted ages. We find that the outliers with SFR M⊙ yr-1 are indeed found to be such a population, they typically located around and with mass-weighted ages –1.5 Gyr, exhibit a higher contribution from older stellar light.
Therefore, both AGN activity and older stellar populations can cause galaxies to deviate from the canonical –SFR relation. AGN-dominated SMGs tend to be younger and more luminous, whereas systems with significant old stellar populations are generally older and located around . If these outliers are excluded, the median and SFR of the remaining sample are L⊙ and 253 M⊙ yr-1, respectively.
@@ V.4 Dust and Dust fraction
The dust mass estimates for our sample appear systematically higher than those reported in previous studies (cunha2015; Dudzevičiūtė et al., 2020b; Hyun et al., 2023). We attribute these differences to the distinct dust models employed: MAGPHYS utilizes modified blackbody spectra to model warm and cold dust components, whereas our CIGALE implementation adopts the dl2014 model, which accounts for dust heating by both star-forming regions and the diffuse interstellar radiation field. This interpretation is supported by liao2024, who report that CIGALE yields dust masses 1.3–1.5 times greater than MAGPHYS estimates. Furthermore, the inclusion of shorter FIR bands in SED fitting introduces systematic biases. Wavelength coverage near the far-infrared emission peak preferentially samples warmer dust components, thereby reducing sensitivity to colder dust and potentially underestimating total dust mass (tacconi2018). Although this effect will be less at high redshifts (tacconi2018), studies suggest a potential underestimate by a factor of 1.2–2 (birkin2021). While methodological variations exist in both approach and results, they do not materially impact subsequent analyses.
@@ V.5 Dust attenuation
Not only SMGs, but also OFGs and UV-selected galaxies exhibit a similar trend. If this observed trend is not an artifact of sample selection or small-number statistics, we interpret it as evidence for a fundamental shift in the dominant mode of cosmic star formation—from an unobscured to a dust-obscured regime (zavala2021). The underlying physical driver may be the gradual buildup of metallicity in galaxies over cosmic time, which eventually reaches a threshold enabling efficient dust production and obscuration (see Section V.4).
@@ V.5.1 Attenuation and Optical Color
Additionally, SMGs are typically selected based on flux in specific bands, resulting in a heterogeneous population. This diversity is evident in their SEDs and radio properties (see paper IV for the discussions of SED and IR-radio correlation). Attenuation curves can vary significantly between different galaxy types, and even within a single galaxy, they may differ substantially along different lines of sight (uematsu2024). This intrinsic diversity further limits the applicability of a single, universal extinction law to the SMG population.
@@ VI Summary
We apply the “super-deblending” technique (Section II.5, Appendix B) to obtain deblended FIR and submillimeter fluxes, and combine these with panchromatic photometry from optical to radio wavelengths to perform SED fitting using CIGALE (Section IV.2). After applying quality cuts based on SED fitting reliability, our final sample comprises 221 SMGs ( with a deblended S850 of mJy) associated with 186 SCUBA-2 soures. Our main results are summarized as follows:
@@
The median is 3.09 0.07 mag (68% range: 1.69-4.44 mag). Optically bright and faint SMGs have mag and mag, respectively. While shows no strong evolution over , we detect a significant drop in attenuation at –not only for SMGS, but also for OFGs and UV-selected galaxies (Figure 22). This phenominent possibly reflecting a transition from unobscured to dust-obscured star formation modes (zavala2021), potentially driven by metallicity accumulation thresholds.
@@ Acknowledgments
This research has made use of the NASA/IPAC Infrared Science Archive, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of the SVO Filter Profile Service “Carlos Rodrigo”, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00.
Y.A. acknowledges the support from the National Key R&D Program of China (2023YFA1608204), the National Natural Science Foundation of China (NSFC grant 12173089), and the Strategic Priority Research Program of the Chinese Academy of Sciences (grant No.XDB0800301).

### v2

@@ V.2 SFRs and Stellar Mass
The stellar mass distribution of SMGs in the SSA22 field is consistent with previous studies, where median masses typically range from – M⊙ (cunha2015; Zavala et al., 2018; Dudzevičiūtė et al., 2020b; Shim et al., 2022; Hyun et al., 2023). Although the values are similar in magnitude, the median SFR in our SSA22 sample is 0.1–0.2 dex lower than in the literature.
From the AS2UDS sample (Dudzevičiūtė et al., 2020b), we selected sources at with radio or 24 counterparts. Their median stellar mass and star formation rate (SFR) are M⊙ and M⊙ yr-1, respectively. Applying the same redshift cut to the multi-band identified S2TDF sample (Hyun et al., 2023) yields median values of M⊙ for stellar mass and M⊙ yr-1 for SFR. Under identical criteria, we selected a conservative sample at with identifications from VLA or MIPS 24 (Table 5). This sample has a median stellar mass of M⊙ and a median SFR of M⊙ yr-1. While the SFR is consistent with the literature values cited above, the stellar mass is somewhat larger, a discrepancy that can likely be attributed to differences in SED fitting methodologies and results (see Section V). Our radio or 24 identified sample also includes a number of “quenched” systems, yet their SFRs are consistent with those reported in literature catalogs. If we further restrict the selection to sources at with radio or 24 counterparts that lie on or above the star-forming main sequence, the median stellar mass and SFR become M⊙ and M⊙ yr-1, respectively. Under the same conditions, the corresponding median values for the AS2UDS subsample are M⊙ and M⊙ yr-1, and for the S2TDF subsample are M⊙ and M⊙ yr-1. Therefore, we suggest that this may arise from relatively weak constraints in the far-infrared during SED fitting, which could lead to greater uncertainty in derived properties and consequently their misclassification.
@@
The median deblended 850 fluxes for counterparts of newly detected sources ( 2.8-3.5 mJy) and S2CLS-detected sources ( 3.5–4.2 mJy) are 3.2 mJy and 4.9 mJy, respectively. These two SMG groups show median SFR and values of 117 M⊙ yr-1, 1.59 1011 M⊙ and 186 M⊙ yr-1, 1.55 1011 M⊙, correspondingly. These results reveal systematic differences in SFR distributions across flux bins, the relation was derived by least-squares fitting: log10(SFR) = (0.49 0.21) log10(S) + (1.78 0.13). When restricting the sample to sources brighter than 4 (or 5) mJy, the median SFR increases to 192 (296) M⊙ yr-1. However, lower SNR sources do not significantly impact the estimated physical properties. For instance, sources with SNR exhibit median SFR and of 175 M⊙ yr-1 and M⊙, while lower-SNR sources (3.5-4) show values of 130 M⊙ yr-1, 1.39 1011 M⊙.
We adopt the main sequence relation from Speagle et al. (2014). We classify the SMGs exhibiting star formation rates three times above this relation as starburst systems, while those with values three times below are categorized as quenched systems (see Paper IV of this series). SMGs located within the main sequence have median SFR and values of 176 M⊙ yr-1 and M⊙, respectively. The ‘starburst system’ exhibit higher SFR but lower galaxy masses, with median values of SFR 510 M⊙ yr-1 and M⊙. In contrast, the ‘quench system’ are on average more massive, with a median of M⊙ — approximately twice the average of full sample — but a median SFR of only 28 M⊙ yr-1.
@@ V.2.1 Mass-to-light Ratio
At typical redshifts for SMGs, the IRAC 5.8 band corresponds approximately to the rest-frame H-band, which serves as a reliable tracer of the dominant stellar population—low-mass stars (Dudzevičiūtė et al., 2020b). As shown in Figure 20, the 151 SMGs with 5.8 detections follow a distribution and trend consistent with the AS2UDS sample: brighter 5.8 emission corresponds to higher stellar mass. A linear fit yields: [M [mag] + (14.97 1.16).
@@ V.3 Infrared Luminosity
The median infrared luminosity (, integrated over 8–1000 in the rest frame) of the sample is L⊙, with a 1 range of – L⊙, and the most luminous source reaches L⊙ (lower-left panel of Figure 19). While slightly lower than literature values, the infrared luminosities of SSA22 SMGs remain comparable in order of magnitude to previous studies.
Similar to the SFR results discussed earlier, when we conservatively restrict our sample to sources with 1 that have radio or 24 counterparts, the median is 3.24 L⊙. For comparison, selecting AS2UDS sources that have multiwavelength counterparts and satisfy 1 yields a median infrared luminosity of 2.88 L⊙. Applying the same redshift cut to the S2TDF sample gives a median value of 3.63 L⊙, in excellent agreement with our result.
@@
Due to the tight correlation between infrared luminosity and star formation activity, exhibits a significant evolution with the cosmic star formation history in Figure 19. Dudzevičiūtė et al. (2020b) found that over the redshift range –4, scales approximately as , but they argue this trend may be biased by the dependence of SED fitting on FIR data. Nevertheless, it is evident that the most luminous SMGs show a clear decline in brightness: HLIRGs are predominantly found at , and at , SMG infrared luminosities typically fall below L⊙.
@@
However, we note that sources residing in the large-scale structure, and those with significant AGN contributions do not systematically deviate from the infrared–SFR relation. We mark two confirmed X-ray AGNs (AGN7 and the quasar ADF22.A1; Lehmer et al., 2009b; Alexander et al., 2016; Monson et al., 2023; Umehata et al., 2025b), yet both lie well within the linear range of the relation between infrared luminosity and SFR. This is consistent with the findings of Monson et al. (2023), who concluded that the majority of AGN-hosting SMGs are indistinguishable from normal star-forming galaxies.
We also identify a group of counterparts with SFRs around 10 M⊙ yr-1 that lie significantly below the relation. Most sources at 1 do not show strong deviations; instead, the outliers are predominantly quiescent identifications. These typically exhibit mass-weighted stellar ages of 1–1.5 Gyr and a high fractional contribution from old stellar populations to the infrared luminosity. As noted by cunha2015; Dudzevičiūtė et al. (2020b), MAGPHYS allows heating of dust by older stellar populations, which can produce excess infrared emission—particularly in galaxies with high mass-weighted ages. This effect may contribute to the apparent offset of these quiescent systems from the canonical IR–SFR relation.
Finally, we find that sources with multiple counterparts do not systematically lie outside the expected relation. Given the low multiplicity fraction in our sample, we conclude that flux misallocation due to source blending does not significantly bias our results.
Therefore, we believe that both AGN activity and older stellar populations can cause galaxies to deviate from the canonical –SFR relation. AGN-dominated SMGs tend to be younger and more luminous, whereas systems with significant old stellar populations are generally older and located around . If these outliers are excluded, the median and SFR of the remaining sample are L⊙ and 253 M⊙ yr-1, respectively.
@@ V.4 Dust and Dust fraction
The dust mass estimates for our sample appear systematically higher than those reported in previous studies (cunha2015; Dudzevičiūtė et al., 2020b; Hyun et al., 2023). We attribute these differences to the distinct dust models employed: MAGPHYS utilizes modified blackbody spectra to model warm and cold dust components, whereas our CIGALE implementation adopts the dl2014 model, which accounts for dust heating by both star-forming regions and the diffuse interstellar radiation field. This interpretation is supported by liao2024, who report that CIGALE yields dust masses 1.3–1.5 times greater than MAGPHYS estimates. Furthermore, the inclusion of shorter FIR bands in SED fitting introduces systematic biases. Wavelength coverage near the far-infrared emission peak preferentially samples warmer dust components, thereby reducing sensitivity to colder dust and potentially underestimating total dust mass (tacconi2018). Although this effect will be less at high redshifts (tacconi2018), studies suggest a potential underestimate by a factor of 1.2–2 (birkin2021). Additionally, contamination during the identification process and potential misallocation of flux may also introduce uncertainties (Section III.3). While methodological variations exist in both approach and results, they do not materially impact subsequent analyses.
@@ V.5 Dust attenuation
The median values for all VLA counterparts, 24 counterparts, 8 counterparts, and color-selected counterparts are 3.10, 2.65, 3.06, and 3.13 mag, respectively. Since most 8 and color-selected counterparts are also identified in the VLA or 24 samples—and radio-detected sources dominate the overall counterpart population—their dust attenuation remains consistent with that of the radio-identified sample. Color-selected galaxies do indeed exhibit redder colors and higher dust attenuation. Considering only those color-selected galaxies without counterparts from other identification methods, the median rises to 3.82 mag. However, this result could be influenced by contamination from lower-redshift sources. When restricted to , the median values for the respective counterpart samples become 3.10, 3.04, 3.08, and 3.09 mag.
@@
Not only SMGs, but also OFGs and UV-selected galaxies exhibit a similar trend. If this observed trend is not an artifact of sample selection or small-number statistics, we interpret it as evidence for a fundamental shift in the dominant mode of cosmic star formation—from an unobscured to a dust-obscured regime (Bourne et al., 2017; Dunlop et al., 2017; Bouwens et al., 2020; zavala2021). The underlying physical driver may be the gradual buildup of metallicity in galaxies over cosmic time, which eventually reaches a threshold enabling efficient dust production and obscuration (see Section V.4).
simpson2014; Dudzevičiūtė et al. (2020b) found that OFGs in the ALESS and AS2UDS samples tend to lie at higher redshifts. Smail et al. (2021) noted that, despite having dust masses and stellar masses comparable to those of other SMGs, these near-infrared–faint SMGs exhibit more compact dust emission and higher star formation rate surface densities. This suggests that such systems may be linked to the formation of compact, spheroidal structures observed in high-redshift galaxies.
@@ V.5.1 Attenuation and Optical Color
Additionally, SMGs are typically selected based on flux in specific bands, resulting in a heterogeneous population (hayward2011; Johnson et al., 2013). For instance, galaxies selected in this way span over two orders of magnitude in both stellar mass and SFR, exhibit a wide range of dust masses and spatial distributions, and encompass diverse evolutionary states. The sample commonly includes both merging or interacting systems and isolated galaxies, with some hosting AGN while others do not. This diversity is evident in their SEDs and radio properties (see Paper IV for the discussions of SED and IR-radio correlation). Attenuation curves can vary significantly between different galaxy types, and even within a single galaxy, they may differ substantially along different lines of sight (uematsu2024). This intrinsic diversity further limits the applicability of a single, universal extinction law to the SMG population.
@@ VI Summary
We apply the “super-deblending” technique (Section II.5, Appendix B) to obtain deblended FIR and submillimeter fluxes, and combine these with panchromatic photometry from optical to radio wavelengths to perform SED fitting using CIGALE (Section IV.2). After applying quality cuts based on SED fitting reliability, our final sample comprises 221 SMGs ( with a deblended S850 of mJy) associated with 186 SCUBA-2 sources. Our main results are summarized as follows:
@@
The median is 3.09 0.07 mag (68% range: 1.69-4.44 mag). Optically bright and faint SMGs have mag and mag, respectively. While shows no strong evolution over , we detect a significant drop in attenuation at –not only for SMGs, but also for OFGs and UV-selected galaxies (Figure 24). This phenomenon possibly reflecting a transition from unobscured to dust-obscured star formation modes (Bourne et al., 2017; Dunlop et al., 2017; Bouwens et al., 2020; zavala2021), potentially driven by metallicity accumulation thresholds.
@@ Acknowledgments
We sincerely thank Ian Smail for his constructive comments and insightful suggestions, which significantly improved the rigor and clarity of this manuscript.
Y.A. acknowledges the support from the National Key R&D Program of China (2023YFA1608204), the National Natural Science Foundation of China (NSFC grant 12173089), the China Manned Space Program with grant no. CMS-CSST-2025-A10, and the Strategic Priority Research Program of the Chinese Academy of Sciences (grant No.XDB0800301).
@@
This work is based on observations taken with the JCMT under project codes M15AI91 and MJLSC02. The MJLSC02 data are part of the SCUBA-2 Cosmology Legacy Survey. This research has made use of the NASA/IPAC Infrared Science Archive, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of the SVO Filter Profile Service “Carlos Rodrigo”, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00.